OpenStudy (anonymous):

Add and write in simplest form:

OpenStudy (anonymous):

\[4y \over y+2 \] + \[8 \over y+2\]

OpenStudy (anonymous):

Answer choices: A. 4 B. \[4y+8 \over y+2 \] C. \[2y+4 \over y\]

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

A

OpenStudy (lalaly):

\[=\frac{4y+8}{y+2} = \frac{4(y+2)}{y+2} = 4\]

OpenStudy (anonymous):

ow ya. should be A